A triangle has side c=2 and angles A=pi/4 and B=pi/6. Find the length of the side opposite A.
1
Expert's answer
2012-01-19T09:57:26-0500
sum of angles pi so angle C=pi-pi/4-pi/6=7pi/12
then we use sine theorem
a/sin(A)=c/sin(C) so a=c*sin(A)/sin(C)=sqrt(2)/sin(7pi/12)